# Learning from galactic rotation curves: a neural network approach

### v1

@@ IV.2 Employing heteroscedastic loss function
Another way to account for aleatoric uncertainty involves changing the loss function used for training the neural network. Consider the case where, for every output corresponding to the parameter prediction, one assigns an additional output that represents the Gaussian variance in that output. This implies that instead of learning a point estimate of the parameter, the neural network now learns the mean and variance of the Gaussian distribution to which the parameter belongs. Hence, in the output layer, for every $k$th parameter in ${\bf P}$, we assign an additional output to be its variance $\sigma_{k}^{2}$. For our neural network, we shall now have an output layer $j=3$ with $N_{j}=10$ outputs, with the first five outputs corresponding to the parameter vector ${\bf\hat{P}}$ and the last five to the uncertainties of the parameters.
@@
$\mathcal{L}_{HS}=\frac{1}{n}\sum_{i=1}^{n}\left[\frac{1}{2p}\sum_{k=1}^{p}\left(e^{-s_{ik}}(y_{ik}-\hat{y}_{ik})^{2}+s_{ik}\right)\right]\ ,$ (15)
where $y_{ik}$ is the $k$th parameter for the $i$th sample in the training set, while $\hat{y}_{ik}\equiv f_{k}({\bf x}_{i},{\bm{\Omega}})$ is the neural network prediction for the same. This is called heteroscedastic loss [48].
@@
IV.2.1 Training and inference using heteroscedastic loss
Recent work [15] has suggested that cosmological parameters inferred from a neural network trained using heteroscedastic loss can infer uncertainties that are similar to those obtained using MCMC. We look at the UGC 5721 galaxy as an example. Utilizing the same architecture and the hyperparameter values as the previous case with noisy input data, we train the neural network for $250$ epochs. We find that the validation loss has converged by this epoch.
Here, to test on unseen data, since the neural networks also predicts uncertainties along with the point estimates of the parameters, we employ a different method than the previous. For $5000$ samples from test data, we plot the difference between the predicted value and target value for each parameter. We also plot $3\sigma_{p}$ for each sample. The plot is shown in Fig. 5, where the pink lines are $3\sigma_{p}$ values while the blue dots are differences between predicted and target parameters. Similar to [15] all samples and parameters, the difference between predicted and target values lies within three times the uncertainty values. Hence, these uncertainties appear to capture the ability of the neural network (with the chosen hyperparameters) to learn the parameters from the training data.
Figure 5: Performance of the neural network for UGC 5721 trained using a heteroscedastic loss function on test data. The blue dots are the difference between inferred and target parameter values, while the pink lines correspond to three times the inferred uncertainty of prediction.
Note that larger differences in $r_{t}$, $r_{s}$ and $\Upsilon_{*}$ and correspondingly larger uncertainties. The larger uncertainties in the $r_{s}$ and $\Upsilon_{*}$ are expected, since rotation curves of dark matter dominated dwarf galaxies are not sensitive to these parameters. For the case of $r_{s}$, since we force $r_{t}\geq 1\ \text{kpc}$ and rotation curves extend only till $\mathcal{O}(5)\ \text{kpc}$ for most galaxies in our sample, a change in the NFW scale radius does not alter the rotation curve significantly. Similarly, because dark matter dominated galaxies have small baryonic contribution to begin with, changing $\Upsilon_{*}$ does not affect the rotation curve much. We shall see in Section V, that this quantification of the uncertainty agrees well with the uncertainty in parameters obtained from MCMC methods.
@@ IV.2.1 Training and inference using heteroscedastic loss
The rotation curve corresponding to the above inferred parameters agrees well with observations for UGC 5721 which can be seen in Fig. 6 by the green curve.
Figure 6: Rotation curves corresponding to the predicted parameters for the neural network trained for UGC 5271 on noisy input data with: (a) MSE loss and mean of multiple parameter predictions (orange) and (b) heteroscedastic loss (green).
We reiterate that only the central value of observed rotation curve is given to the neural network during inference, hence, there is only one prediction for the parameters and their uncertainties. Then network has learned the uncertainty implicitly from the noisy training data.
@@
Figure 7: Rotation curves corresponding to the predicted parameters for both the cases studied: (a) MSE loss function with noisy input data with multiple realizations of the observed rotation curves from Section IV.1 (orange curve) , and (b) Heteroscedastic loss function with noisy inputs from Section IV.2 (green curve). The reduced $\chi^{2}$ are also shown for each case.
The story remains the same for all galaxies in the samples, rotation curves for which are shown in Fig. 7. Notice the larger uncertainties associated with $r_{s}$ and $\Upsilon_{*}$ for the other galaxies as well, which is remarkably consistent with the expectation that the rotation curves for DM dominated galaxies are not very sensitive to $r_{s}$ and $\Upsilon_{*}$ values. This implies that uncertainty in the heteroscedastic loss function captures in some sense how much the effect a parameter will have on a rotation curve. This is interesting since the neural network has only approximated the functional dependence between parameters and rotation curves by looking at various examples.
@@
Galaxy Predictions
DDO 154 $1.95_{-0.3}^{+0.3}$ $5350.71_{-292.16}^{+292.16}$ $3.50_{-0.87}^{+0.87}$ $6.87_{-3.90}^{+3.90}$ $0.61_{-0.12}^{+0.12}$
ESO 444-G084 $5.26_{-1.07}^{+1.07}$ $5055.19_{-363.85}^{+363.85}$ $1.26_{-0.27}^{+0.27}$ $9.53_{-3.38}^{+3.38}$ $0.56_{-0.14}^{+0.14}$
UGC 5721 $2.21_{-0.34}^{+0.34}$ $3052.34_{-163.26}^{+163.26}$ $2.55_{-0.72}^{+0.72}$ $7.12{-4.04}^{+4.04}$ $0.66_{-0.13}^{+0.13}$
UGC 5764 $3.87_{-0.43}^{+0.43}$ $4769.11_{-192.48}^{+192.48}$ $2.00{-0.51}^{+0.51}$ $4.75_{-3.74}^{+3.74}$ $0.57_{-0.14}^{+0.14}$
UGC 7524 $1.55_{-0.52}^{+0.52}$ $4370.11_{-422.96}^{+422.96}$ $3.02_{-1.43}^{+1.43}$ $9.56{-2.86}^{+2.86}$ $0.57_{-0.14}^{+0.14}$
UGC 7603 $2.61_{-0.65}^{+0.65}$ $4243.85_{-368.33}^{+368.33}$ $2.2-_{-0.77}^{+0.77}$ $7.43_{-4.04}^{+4.04}$ $0.56_{-0.14}^{+0.14}$
UGCA 444 $6.84_{-1.61}^{+1.61}$ $7958.19_{-742.22}^{+742.22}$ $1.47_{-0.35}^{0.35}$ $8.23_{-3.91}^{+3.91}$ $0.57_{-0.14}^{+0.14}$
Table 3: Parameters and their uncertainties obtained by feeding observed rotation curves as input to neural networks trained using a heteroscedastic loss function.
V Comparison with MCMC
@@
Recall from Section IV.1, that for a fixed galaxy, we sample from the Gaussian $\mathcal{N}(v_{obs}(r),\sigma_{r}^{2})$ for all radius values to generate multiple realizations of the observed rotation curve. One can think of this as sampling curves from the vicinity of the observed rotation curve while being consistent with observations. For a sufficiently accurate neural network, one can then think of the multiple parameter predictions (using the multiple realizations as input) as a ‘chain’ of parameters that satisfy the observed rotation curve of that galaxy. One can then plot 2D and 1D projections of the parameter values for this chain in a corner plot to visualize how the parameters are distributed.
@@
Therefore, we proceed to use the ‘chain’ of $1000$ inferred parameters, to obtain the median and $1\sigma$ confidence interval for each parameter for every galaxy in our sample as shown by the orange points with error bars in Fig. 9. One can also use the chain of parameters to plot contours of the 2D joint and 1D marginalized distributions in a corner plot. Contours obtained for the case of UGC 5721, as an example, are shown in red in Fig. 8.
To compare with Bayesian inference, we sample the posterior using an ensemble sampler incorporated in the emcee [^2] https://emcee.readthedocs.io/en/v2.2.1/ python package, for all galaxies in the sample. We use the same model described in Section II.1, where the theory rotation curve is parameterized by $5$ parameters in Eq. (7). We use the same uniform priors as those defined in Table 1. Here, the log-likelihood is written as
$\mathrm{ln}\mathcal{L}_{\text{MCMC}}=-\frac{1}{2}\sum_{i}\left(\frac{v_{obs}(r_{i})-v_{th}(r_{i};{\bf P})}{\sigma_{i}}\right)^{2}\ .$ (16)
We then run the emcee sampler with 48 walkers (at different random initial points in the parameter space) for $20,000$ steps, out of which the initial $10,000$ are removed to account for the burn-in period. The median and $1\sigma$ confidence intervals are shown in Fig. 9 by the yellow points with error bars for all parameters for every galaxy in our sample. These chains can also be used to plot $2$D and $1$D posterior distributions, which are shown for the case of UGC 5721 in Fig. 8 by the green contours.
Figure 8: $2$D and $1$D distributions of parameters obtained using our approach as described in Section IV.1 (in red), plotted over joint and marginal distributions obtained using MCMC for UGC 5721 (in green). The vertical dashed lines (red and green) in the $1$D histograms are the $50\%$ quantile values for both approaches. The $2$D contours show $\sim 39.3\%$ and $86.4\%$ confidence regions.
Note that, for the neural network trained using the heteroscedastic loss function, one does not get multiple parameter estimates. Here, the uncertainty is learned implicitly during training, and hence one cannot plot the distribution. We therefore plot the point-estimate and uncertainties obtained using this method in Fig. 9 by the green points with error bars.
@@
  • The median values obtained from all three approaches are close, i.e., the median values ($50\%$ quantile) using our approach lies within the $1\sigma$ confidence interval obtained from the MCMC approach.
  • From Fig. 8, it is also clear that contours obtained using the ANN approach seem to capture correct correlations between some parameters, for instance, the positive correlation between $m$ and $s$, the negative correlation between $r_{t}$ and $m$ as well as $r_{t}$ and $s$.
  • Predicted uncertainties are also similar for both methods for $m$, $s$ and $r_{t}$. However, the neural networks severely under-estimates the uncertainties for $r_{s}$ and $\Upsilon_{*}$, which, as we discussed in Section IV.1, do not affect the rotation curve significantly. This is captured correctly by the by MCMC method, which gives large uncertainties for $r_{s}$ and $\Upsilon_{*}$ for all galaxies in Fig. 8.
VI Discussion and conclusion
@@
In this work, unlike a likelihood-based approach, we train neural networks with rotation curves as input and the model parameters as output using a large sample of simulated rotation curves whose parameter values are already known. By looking at samples in the training data, the neural network updates its internal adjustable parameters (called weights and biases) to approximate a function $f:\mathbb{R}^{N_{obs}}\rightarrow\mathbb{R}^{P}$ where $N_{obs}$ is the number of observed velocities in a rotation curve and $P=5$ is the number of parameters, which are: ULDM particle mass $m$, the scale parameter $s$ associated with the mass of the core, transition radius $r_{t}$, i.e. the radius at which the core profile transitions to a NFW profile, the scale radius of the NFW profile $r_{s}$, and $\Upsilon_{*}$, the stellar mass-to-light ratio which tunes the baryonic contribution to the rotation curves (See Section II.1 for a detailed discussion on the model and dataset used).
(a) ULDM mass $m\ (\text{eV})$
(b) Scale parameter $s$
(c) Transition radius $r_{t}\ (\text{kpc})$
(d) Scale radius $r_{s}\ (\text{kpc})$
(e) Stellar mass-to-light ratio $\Upsilon_{*}\ (M_{\odot}/L_{\odot})$
Figure 9: Inferred parameters and their uncertainties using multiple realization method from Section IV.1 (orange), heteroscedastic loss function from Section IV.2 (green) and MCMC runs using emcee in Section V (yellow).
@@
    Neural networks trained on noisy simulated rotation curves perform much better than the noiseless case when confronted with observed rotation curves. The noise included during training appears to capture the fact that observed rotation curves are not “smooth” functions (see Fig. 4 and the discussion around it).
@@
    To quantify the uncertainty associated with the inferred parameters based on the uncertainty in the observations, we generated multiple realizations of the observations as described in Section IV.1, which lead to multiple parameter estimations using observations, resembling a chain of parameters. The median values obtained using this method agree well lead to rotation curves which agree well observed rotation curves, as shown in Fig. 7 by the orange curves.
@@
    Instead of mean-squared-error (MSE) loss function, using a heteroscedastic loss function in eq. (15) to train the network enables us to learn uncertainty during training itself. When confronted with observed rotation curves, the uncertainties for $r_{s}$ and $\Upsilon_{*}$ are high for all galaxies. Though this result is expected, it is not captured by the uncertainties obtained from the multiple realizations method (see Section IV.2.1). Furthermore, the point-estimates of the parameters also fit the observations just as well as the MSE case as seen by the green curves in Fig. 7.
@@
    Comparison with MCMC in Section V suggests that while the multiple realizations method can capture some of the correlations between a few parameters, namely $m$, $s$ and $r_{t}$, it struggles when confronted with hard-to-constrain parameters like $r_{s}$ and $\Upsilon_{*}$ which don’t affect the rotation curves in our sample significantly.
@@
    For the case of neural networks trained using the heteroscedastic loss function, while one cannot obtain $2$D and $1$D distribution of parameters, the ANNs are able to capture the amount of uncertainty better for the hard-to-constrain parameters ($r_{s}$ and $\Upsilon_{*}$). Indeed, as shown in Fig. 9, for almost all galaxies, the parameters and the uncertainties inferred by neural networks with the heteroscedastic loss function agree well with those obtained from MCMC analysis.
@@
In summary, we have shown that with the chosen architecture and hyperparameters, neural networks can prove to be a useful tool in obtaining parameter estimates that can describe observed rotation curves well, i.e. with a small $\chi^{2}_{red}$. Our work, along with other recent works [10, 11, 15, 20, 21] demonstrates that, the use of neural networks for this class of problems can be a useful complementary approach to standard likelihood-based approaches. In particular, our work agrees well with the recent results obtained in [15], where model parameters as well as their uncertainties obtained using an ANN trained with a heteroscedastic loss function agreed well with those obtained using MCMC.
@@
Before closing we would like to note some caveats and future prospects: (a) The true values of each parameter are assumed to be independently and normally distributed in the heteroscedastic loss function, which may not be the case. (b) We have not carried out a full hyperparameter exploration to obtain an optimal neural network. It is likely that there exists a better choice of hyperparameters that perform better by giving better estimates of the parameter values and the associated uncertainties for this dataset. (c) Information about the radius values for which velocities (observed or simulated) are obtained has not been used in this analysis, since the input vector is just a list of velocities. Therefore, the neural networks we have trained cannot differentiate between two rotation curves with the same velocities but different observed radii. An interesting direction would be to train a neural network capable of handling input rotation curves from multiple galaxies with a varied range of radii and velocities to infer a single parameter value for the fundamental parameter of our model, i.e., ULDM particle mass $m$, while inferring different galaxy specific parameters. We leave this exploration to future work. (d) In this work, we deal with the vanilla fuzzy dark matter model, i.e., a scalar field with negligible self-interactions. In our recent work [37] we have demonstrated that repulsive self-interactions can satisfy observed rotation curves while also satisfying an empirical core-halo mass relation, unlike the case of no self-interactions [50]. It could then be interesting to include the effect of $\lambda\varphi^{4}$ kind of self-interactions for the ultralight scalar field, and infer $\lambda$ along with the parameters we considered in this work.
Acknowledgements.

### v2

@@ 4.2 Employing heteroscedastic loss function
Another way to account for aleatoric uncertainty involves changing the loss function used for training the neural network. Consider the case where, for every output corresponding to the parameter prediction, one assigns an additional output that represents the Gaussian variance in that output. This implies that instead of learning a point estimate of the parameter, the neural network now learns the mean and variance of the Gaussian distribution to which the parameter belongs. Hence, in the output layer, for every $k$th parameter in $\bf P$}, we assign an additional output to be its variance $\sigma_{k}^{2}$. For our neural network, we shall now have an output layer $j=3$ with $N_{3}=10$ outputs, with the first five outputs corresponding to the parameter vector $^{\bf\hat{P}}$ and the last five to the uncertainties of the parameters.
@@
$\mathcal{L}_{HS}=\frac{1}{n}\sum_{i=1}^{n}\left[\frac{1}{2p}\sum_{k=1}^{p}\left(e^{-s_{ik}}(y_{ik}-\hat{y}_{ik})^{2}+s_{ik}\right)\right]\ ,$ (4.1)
where $y_{ik}$ is the $k$th parameter for the $i$th sample in the training set, while $\hat{y}_{ik}\equiv f_{k}({\bf x}_{i},{\boldsymbol{\Omega}})$ is the neural network prediction for the same. This is called heteroscedastic loss [54]. Similar loss functions have been utilized in parameter estimation recently in the context of cosmological parameters [15, 55] as well as gravitational wave parameters [56].
@@
4.2.1 Training and inference using heteroscedastic loss
As with the previous section, we shall first look at the UGC 5721 galaxy as an example. Utilizing the same architecture and the hyperparameter values as the previous case with noisy input data, we train the neural network for $250$ epochs. We find that the validation loss has converged by this epoch.
Here, to test on unseen data, since the neural networks also predict uncertainties along with the point estimates of the parameters, we employ a different method than earlier. For $500$ samples from test data, we plot the difference between the predicted value and target value for each parameter. We also plot $3\hat{\sigma}_{p}$ for each sample. The plot is shown in figure 5, where the pink lines are $3\hat{\sigma}_{p}$ values while the red dots are differences between predicted and target parameters. Similar to [15], for most samples, the difference between predicted and target value of each parameter lies within three times the uncertainty values. Hence, these uncertainties appear to capture the ability of the neural network (with the chosen hyperparameters) to learn the parameters from the training data.
Refer to caption Refer to caption Refer to caption
(a) ULDM mass, $m$ (b) scale parameter, $s$ (c) transition radius, $r_{t}$
Refer to caption Refer to caption
(d) scale radius, $r_{s}$ (e) stellar mass-to-light ratio, $\Upsilon_{*}$
Figure 5: Performance of the neural network for UGC 5721 trained using a heteroscedastic loss function on test data. The red dots are the difference between inferred and target parameter values, e.g. $\Delta m=m_{\text{true}}-m_{\text{inferred}}$, while the pink lines correspond to three times the inferred uncertainty of prediction.
Note that larger differences in $r_{t}$, $r_{s}$ and $\Upsilon_{*}$ and correspondingly larger uncertainties. The larger uncertainties in the $r_{s}$ and $\Upsilon_{*}$ are expected, since rotation curves of dark matter dominated dwarf galaxies are not sensitive to these parameters. For the case of $r_{s}$, since the inner regions are described by a soliton core, i.e., we force $r_{t}\geq 1\ \text{kpc}$ and rotation curves extend only till $\mathcal{O}(5)\ \text{kpc}$ for most galaxies in our sample, a change in the NFW scale radius does not alter the rotation curve significantly. Similarly, because dark matter dominated galaxies have small baryonic contribution to begin with, changing $\Upsilon_{*}$ does not affect the rotation curve much. We shall see in section 5, how this quantification of the uncertainty compares with the uncertainty in parameters obtained from MCMC methods.
@@ 4.2.1 Training and inference using heteroscedastic loss
Figure 6: Rotation curves corresponding to the predicted parameters for the neural network trained for UGC 5271 on noisy input data with: (a) MSE loss and mean of multiple parameter predictions (purple dashed) and (b) heteroscedastic loss (orange).
The rotation curve corresponding to the above inferred parameters agrees well with observations for UGC 5721 which can be seen in figure 6 by the orange curve. We reiterate that only the central value of observed rotation curve is given to the neural network during inference, hence, there is only one prediction for the parameters and their uncertainties. Then network has learned the uncertainty implicitly from the noisy training data.
The story remains the same for all galaxies in the samples, rotation curves for which are shown in figure 7, while the parameter values and their uncertainties are listed in Table 2. Notice the larger uncertainties associated with $r_{s}$ and $\Upsilon_{*}$ for the other galaxies as well, which is remarkably consistent with the expectation that the rotation curves for DM dominated galaxies with $r_{t}\geq 1\ \text{kpc}$ are not very sensitive to $r_{s}$ and $\Upsilon_{*}$ values. This implies that uncertainty in the heteroscedastic loss function captures, in some sense, the effect a parameter will have on a rotation curve. This is interesting since the neural network has only approximated the functional dependence between parameters and rotation curves by looking at various examples.
@@
Figure 7: Rotation curves corresponding to the predicted parameters for each of the three cases studied:(a) MSE loss function with noisy input data with multiple realizations of the observed rotation curves from section 4.1 (purple dashed curves) , and (b) Heteroscedastic loss function with noisy inputs from section 4.2 (orange curves). The reduced $\chi^{2}$ are also shown for each case.
@@
5 Comparison with MCMC
@@
Recall from section 4.1, that for a fixed galaxy, we sample from the Gaussian $\mathcal{N}(v_{obs}(r),\sigma_{r}^{2})$ for all radius values to generate multiple realizations of the observed rotation curve. One can think of this as sampling curves from the vicinity of the observed rotation curve while being consistent with observations. For a sufficiently accurate neural network, one can then think of the multiple parameter predictions (using the multiple realizations as input) as a ‘chain’ of parameters that satisfy the observed rotation curve of that galaxy. One can then plot 2D and 1D projections of the parameter values for this chain in a corner plot to visualize how the parameters are distributed.
@@
Therefore, we proceed to use the ‘chain’ of $1000$ inferred parameters, to obtain the median and $1\sigma$ confidence interval for each parameter for every galaxy in our sample as shown by the purple points with error bars in figure 9. One can also use the chain of parameters to plot contours of the 2D joint and 1D marginalized distributions in a corner plot. Contours obtained for the case of UGC 5721, as an example, are shown in red in figure 8.
To compare with Bayesian inference, we sample the posterior using an ensemble sampler incorporated in the emcee [^2] https://emcee.readthedocs.io/en/v2.2.1/ python package, for all galaxies in the sample. We use the same model described in section 2.1, where the theory rotation curve is parameterized by $5$ parameters in eq. (2.7). We also use the same uniform priors as those defined in Table 1. Here, the log-likelihood is written as
$\mathrm{ln}\mathcal{L}_{\text{MCMC}}=-\frac{1}{2}\sum_{i}\left(\frac{v_{obs}(r_{i})-v_{th}(r_{i};{\bf P})}{\sigma_{i}}\right)^{2}\ .$ (5.1)
We then run the emcee sampler with 48 walkers (at different random initial points in the parameter space) for $20,000$ steps, out of which the initial $10,000$ are removed to account for the burn-in period. The median and $1\sigma$ confidence intervals are shown in figure 9 by the green points with error bars for all parameters for every galaxy in our sample (they are also listed in Table 2). These chains can also be used to plot 22D and 11D posterior distributions, which are shown for the case of UGC 5721 in figure 8 by the green contours.
Figure 8: 22D and 11D distributions of parameters obtained using our approach as described in section 4.1 (in red), plotted over joint and marginal distributions obtained using MCMC for UGC 5721 (in green). The vertical dashed lines (red and green) in the 11D histograms are the $50\%$ quantile values for both approaches. The 22D contours show $\sim 39.3\%$ and $86.4\%$ confidence regions.
Note that, for the neural network trained using the heteroscedastic loss function, one does not get multiple parameter estimates. Here, the uncertainty is learned implicitly during training, and hence one cannot plot the joint distributions. We therefore plot the point-estimate and uncertainties obtained using this method in figure 9, denoted by the orange points with error bars.
@@
  • The median values obtained from the multiple realizations approach as well as the point-estimates obtained from neural networks trained using heteroscedastic loss function lie within $1\sigma$ intervals obtained from the MCMC approach for most parameters and galaxies.
  • As seen from the case of UGC 5721 in figure 8, it is also clear that contours obtained using the ANN approach seem to capture correct correlations between some parameters, for instance, the positive correlation between $m$ and $s$, the negative correlation between $r_{t}$ and $m$ as well as $r_{t}$ and $s$.
  • Uncertainties obtained from all three methods for $m$, $s$ and $r_{t}$ are also similar. However, the multiple realizations approach severely under-estimates the uncertainties for $r_{s}$ and $\Upsilon_{*}$, which, as we discussed in section 4.2, do not affect the rotation curves significantly. This is captured correctly by the neural networks trained using heteroscedastic loss, which gives large uncertainties for $r_{s}$ and $\Upsilon_{*}$ for all galaxies in figure 9, agreeing well with the MCMC approach.
5.1 Uncertainties obtained using heteroscedastic loss function
At this stage it is important to note that, one must be cautious in the interpretation of uncertainties in the field of machine learning and in parameter estimation problems from cosmology and astrophysics [57]. For instance, the term ‘uncertainty’ is used to describe the $\sigma_{ik}$ value obtained from eq. (4.1) predicted by an ANN using the heteroscedastic loss function, as well as the $1\sigma$ value obtained from an MCMC chain. These two uncertainties may not be equivalent. This can be understood by noting that in the MCMC approach, the uncertainty propagates from the data to the parameter posteriors via the likelihood function, given a distribution for the errors associated with observations (in this case, Gaussian). On the other hand, the uncertainties predicted using the heteroscedastic loss function characterize, to an extent, the difference between predicted and target parameter values, based on the input simulated rotation curves.
It is therefore interesting to note that the uncertainties obtained using both approaches are similar, as seen in figure 9; in particular for the parameters that do not affect rotation curves much and thus have larger uncertainties, i.e., NFW scale radius $r_{s}$ and stellar mass-to-light ratio $\Upsilon_{*}$. It is worth noting that heteroscedastic loss functions have been used previously to obtain uncertainties associated with predictions [55, 56, 16, 15]. In particular, in [15] the authors utilized this loss function for parameter estimation of cosmological parameters from $H(z)$ data. For this case, the values of the uncertainties obtained using the heteroscedastic loss function were also found to be similar to those obtained using the MCMC approach. This suggests that the heteroscedastic loss function needs to be explored further parameter estimation problems utilizing neural networks in their pipelines.
In this work, while we have focused on exploring an alternative, complementary approach to parameter estimation using neural networks, we leave the exploration of the connection between uncertainties obtained in ANN and MCMC approaches for future work.
6 Discussion and conclusion
@@
In this work, unlike a likelihood-based approach, we train neural networks with rotation curves as input and the model parameters as output using a large sample of simulated rotation curves whose parameter values are already known.  [^3] There are various other approaches that can be used here to carry out parameter estimation along with uncertainty quantification using neural networks including Bayesian neural networks (BNNs) and normalizing flows (see [59, 60, 61, 62] for recent work).. By looking at samples in the training data, the neural network updates its internal adjustable parameters (called weights and biases) to approximate a function $f:\mathbb{R}^{N_{obs}}\rightarrow\mathbb{R}^{P}$ where $N_{obs}$ is the number of observed velocities in a rotation curve and $P=5$ is the number of parameters, which are: ULDM particle mass $m$, the scale parameter $s$ associated with the mass of the core, transition radius $r_{t}$, i.e. the radius at which the core profile transitions to a NFW profile, the scale radius of the NFW profile $r_{s}$, and $\Upsilon_{*}$, the stellar mass-to-light ratio which tunes the baryonic contribution to the rotation curves (See section 2.1 for a detailed discussion on the model and dataset used).
(a) ULDM mass, $m$
Refer to caption Refer to caption
(b) scale, $s$ (d) scale radius, $r_{s}$
Refer to caption Refer to caption
(c) transition radius, $r_{t}$ (e) stellar mass-to-light ratio, $\Upsilon_{*}$
Figure 9: Inferred parameters and their uncertainties using multiple realization method from section 4.1 (purple), heteroscedastic loss function from section 4.2 (orange) and MCMC runs using emcee in section 5 (green).
@@
    Neural networks trained on noisy simulated rotation curves perform much better than the noiseless case when confronted with observed rotation curves. The noise included during training appears to capture the fact that observed rotation curves are not “smooth” functions (see figure 4 and the discussion around it).
@@
    To quantify the uncertainty associated with the inferred parameters based on the uncertainty in the observations, we generated multiple realizations of the observations as described in section 4.1, which lead to multiple parameter estimates, resembling a chain of parameters. The median values obtained using this method lead to rotation curves which agree well with observed rotation curves, as shown in figure 7 by the purple curves.
@@
    However, comparison with parameter estimation using MCMC in section 5 suggests that while the multiple realizations method can capture some of the correlations between a few parameters, namely $m$, $s$ and $r_{t}$, it struggles when confronted with hard-to-constrain parameters like $r_{s}$ and $\Upsilon_{*}$ which don’t affect the rotation curves in our sample significantly (see Figs. 8 and 9).
@@
    Instead of the mean-squared-error (MSE) loss function, using a heteroscedastic loss function in eq. (4.1) to train the network enables us to learn the uncertainty associated with each parameter during training itself. When confronted with observed rotation curves, the ANNs perform just as well as the networks trained using MSE loss function (see figure7).
@@
    For the case of neural networks trained using the heteroscedastic loss function, while one cannot obtain 22D and 11D distribution of parameters, the ANNs predict higher uncertainties for the $r_{s}$ and $\Upsilon_{*}$ parameters compared to the multiple realizations approach. Large uncertainties for $r_{s}$ and $\Upsilon_{*}$ are expected since the rotation curves are not very sensitive to a change in these parameters, which is also reflected in the large uncertainties obtained using the MCMC approach as shown in figure 9 in green.
@@
Over the last few years, there has been a lot of interest in utilizing neural networks for parameter estimation problems in astrophysics and cosmology (see section 3.2 of [63] for a recent summary). Many of these methods utilize neural networks as a part of a larger likelihood-free inference pipeline [12, 11, 56, 15, 62, 21, 22]. In this work, we have explored a simple scenario where neural networks are trained to directly output parameter estimates given rotation curves as input, while also comparing different methods of quantifying uncertainties associated with each parameter to those obtained using the traditional Bayesian approach. We have found that, given the chosen architecture and hyperparameters, neural networks can prove to be a useful tool in obtaining parameter estimates that can describe observed rotation curves well, i.e. with a small $\chi^{2}_{red}$. Our work, along with other recent work [10, 11, 15, 21, 22] demonstrates that, the use of neural networks for this class of problems can be a useful complementary approach to standard likelihood-based approaches. However, we would like to note that, (a) uncertainties obtained using the heteroscedastic loss function may not be equivalent to the familiar Bayesian posterior (as we have discussed in section 5.1), and (b) the total time taken for the generation of training samples and training the neural network is much longer than the time taken to run an MCMC sampler. This implies that the ANN approach that we have incorporated in this work cannot surpass or replace the traditional MCMC approach yet. It is also worth noting that, the simple neural network approach we have considered in section 4.2.1 agrees well with the recent results obtained in [15], where the authors found that model parameters as well as their uncertainties obtained using an ANN trained with a heteroscedastic loss function compared well with the MCMC approach.
@@
Before closing we would like to note some caveats and future prospects: (a) The true values of each parameter are assumed to be independently and normally distributed in the heteroscedastic loss function, which may not be the case. (b) We have not carried out a full hyperparameter exploration to obtain an optimal neural network. It is likely that there exists a better choice of hyperparameters that perform better by giving better estimates of the parameter values and the associated uncertainties for this dataset as well as decrease the total amount of time taken during the training phase. (c) Information about the radius values for which velocities (observed or simulated) are obtained has not been used in this analysis, since the input vector is just a list of velocities. Therefore, the neural networks we have trained cannot differentiate between two rotation curves with the same velocities but different observed radii. An interesting direction would be to train a neural network capable of handling input rotation curves from multiple galaxies with a varied range of radii and velocities to infer a single parameter value for the fundamental parameter of our model, i.e., ULDM particle mass $m$, while inferring different galaxy specific parameters. We leave this exploration to future work. (d) In this work, we deal with the vanilla fuzzy dark matter model, i.e., a scalar field with negligible self-interactions. In our recent work [39] we have demonstrated that repulsive self-interactions can satisfy observed rotation curves while also satisfying an empirical core-halo mass relation, unlike the case of no self-interactions [40]. It could then be interesting to include the effect of $\lambda\varphi^{4}$ kind of self-interactions for the ultralight scalar field, and infer $\lambda$ along with the parameters we considered in this work.
Acknowledgments